# Improved Belief Propagation Decoding algorithms for surface codes

### v2

@@ Abstract
Quantum error correction is crucial for universal fault-tolerant quantum computing. Highly accurate and low-time-complexity decoding algorithms play an indispensable role in making sure quantum error correction works. Among existing decoding algorithms, belief propagation (BP) is notable for its nearly linear time complexity and general applicability to stabilizer codes. However, BP’s decoding accuracy without post-processing is unsatisfactory in most situations. This article focuses on improving the decoding accuracy of BP over GF(4) for surface codes. We first propose Momentum-BP and AdaGrad-BP, inspired by machine learning optimization techniques, to reduce oscillation in message updating and break the symmetric trapping sets. We further propose EWAInit-BP, which adaptively updates initial probabilities and provides a 1 to 3 orders of magnitude improvement over traditional BP for planar surface code, toric code, and XZZX surface code without any post-processing method, showing high decoding accuracy even under parallel scheduling. The theoretical $O(1)$ time complexity under parallel scheduling and high accuracy of EWAInit-BP make it a promising candidate for high-precision real-time decoders. Meanwhile, the ideas of the Momentum-BP, AdaGrad-BP and EWAInit-BP provide promising approaches to improve the decoding accuracy of BP to get rid of its reliance on post-processing.
@@ I Introduction
Information in quantum computing devices is highly susceptible to noises, leading to information loss and computational errors [1]. To address this issue, quantum error correction (QEC) introduces redundant physical qubits and encodes the information in certain subspace of the state space of all the physical qubits to protect the information. So far at least, QEC is an essential step towards universal fault-tolerant quantum computation [2, 3, 4]. To make sure QEC work, the decoder of a QEC code should output the estimation of error fast and accurate according to the error syndromes.
@@
Several decoding algorithms have achieved notable success, including the minimum-weight perfect matching (MWPM) [5] algorithm, union-find (UF) algorithm [6], maximum likelihood (ML) algorithm [7], matrix-product-states-based (MPS-based) algorithm [7], renormalization group algorithm [8], and neural-network-based (NN-based) algorithm [9, 10, 11, 12, 13, 14, 15, 16]. Among these, ML, MWPM, and MPS-based algorithms offer high accuracy but come with high complexity. In contrast, the UF algorithm provides near-linear complexity with slightly lower accuracy. NN-based decoders have to balance scalability, decoding accuracy, and time cost [13], and may struggle with generalizability across different codes. Consequently, existing decoding algorithms rarely meet the requirements for both high accuracy and low complexity in quantum error correction decoding [17, 18].
@@
Belief Propagation (BP) algorithm [19] are applicable to any stabilizer code and can achieve a complexity of $O(j\tau N)$ [20], where $j$ is the average weight of the rows/columns of the check matrix, and $\tau$ is the number of iterations. Additionally, parallel scheduling of BP can be efficiently implemented in hardware, theoretically achieving constant-time complexity [21]. However, traditional BP exhibit low decoding accuracy for many quantum error correcting codes and even cannot achieve the code capacity threshold for surface codes.
@@
BP can be implemented over both GF(4) [22, 23, 24, 25, 26, 20, 27] and GF(2) [28, 29, 30, 31], where the Pauli operators are represented. There have been some efforts improving decoding accuracy of BP over GF(4) and GF(2). Over GF(4), MBP [20] achieves a threshold of 14.5% to 16% of the planar surface code by modifying the step size in the a posteriori probability updating. However, MBP is limited to serial scheduling, and the outer loop for searching step sizes increases its time complexity. Over GF(2), GBP [31] achieves a 14% threshold by combining generalized belief propagation with multiple rounds of decoding. This algorithm can be implemented in parallel, but the number of outer loops increases with the code length, adding to its time complexity. BP-OTS [32] periodically detects and breaks trapping sets using manually set parameters, allowing parallel implementation while maintaining the original complexity. It outperforms MBP for short codes but does not achieve a significant threshold.
@@
This article employs three strategies to improve BP: optimizing message updates by optimization methods in machine learning, and optimizing a priori probabilities by utilizing a posteriori probabilities output by previous iterations. First, inspired by the similarity between BP message update and gradient descent of the energy function, we leverage optimization methods in machine learning to optimize the message update rules, and proposed Momentum-BP and AdaGrad-BP. In addition to improving decoding accuracy, experiments on trapping sets in surface codes show that Momentum-BP and AdaGrad-BP can smooth the message update process and reduce oscillation amplitude.
Furthermore, realizing the importance of a priori and a posteriori probabilities involved in the iteration of BP, we propose EWAInit-BP which dynamically updates the a priori probabilities involved in each iteration by utilizing the a posteriori probabilities output by previous iterations. Experiments on short toric codes, planar surface codes, and XZZX surface codes (under bias noise) show that, under both parallel scheduling and serial scheduling, EWAInit-BP outperforms existing BP improvements without post-processing and outer loops. Notably, the time complexity of all proposed improvements remains $O(j\tau N)$. Thus, EWAInit-BP can be implemented with theoretical constant complexity in parallel while maintaining decoding accuracy.
@@ II.1 Surface codes
A stabilizer code is defined by an abelian subgroup $\mathcal{S}$ of the Pauli group $\mathcal{P}_{n}$ excluding $-I$, composed of operators $S\in\mathcal{P}_{n}$ such that $S|\psi\rangle=|\psi\rangle$ for all codewords $|\psi\rangle$ in the code space. The code space is the simultaneous $+1$ eigenspace of all elements of $\mathcal{S}$. The stabilizers must commute with each other, which means $[S_{i},S_{j}]=0$ for $\forall S_{i},S_{j}\in\mathcal{S}$. The dimension of the code space is determined by the number of independent generators of $\mathcal{S}$. For a stabilizer code with $N$ physical qubits and $K$ logical qubits, $\mathcal{S}$ has $N-K$ independent generators.
@@
The stabilizers are used to detect errors by measuring their eigenvalues. The set of measurement outcomes, known as the error syndrome, contains some information about the errors on physical qubits. The logical operators of a stabilizer code are Pauli operators that commute with all elements of $\mathcal{S}$ but are not in $\mathcal{S}$, thus in $\mathcal{N(S)\backslash S}$ where $\mathcal{N(S)}=\{E:ES=SE,\ \forall S\in\mathcal{S}\}$. They act on the encoded logical qubits without disturbing the code space. Due to this, the logical operators can be also regarded as undetectable errors, in the case that they are not performed on the physical qubits according to our requirements. The code distance $D$ is defined as the minimum weight of logical operators. A quantum stabilizer code can be represented by $[[N,K,D]]$.
@@
The most eye-catching stabilizer codes in recent years might be surface codes. Surface codes are a class of quantum Low-Density Parity-Check (LDPC) codes [39] where qubits are placed on a 2-dimensional lattice, relying on the topology of the surface to protect quantum information [40]. Logical qubits are encoded in global degrees of freedom, and larger lattices provide larger code distances. There are some common properties of surface codes, such as the code distance scales with the code length as $O(\sqrt{N})$, while the code rate is asymptotically zero. In this article, toric codes refer to the ones with closed boundary, and planar surface codes refer to the ones with open boundary. The stabilizers of these two types are all Pauli $X$ or Pauli $Z$. Besides, XZZX surface codes refer to the toric codes whose stabilizers are in the form of $XZZX$.
@@
A $[[2L^{2},2,L]]$ toric code has a 2-dimensional torus topology, and the closed boundaries makes it translationally invariant to the error syndrome, which is beneficial to decoding algorithms (such as BP and neural networks). The Euler characteristic of the torus is 0, but the X-type and Z-type stabilizers of the toric code both have a redundancy, resulting in a total of two encoded logical qubits.
@@
A $[[2L^{2}-2L+1,1,L]]$ planar surface code is defined on a surface with open boundaries, which makes the stabilizers near the boundaries having lower weights. This reduction in stabilizer weight results in a slight decrease in the error correction capability compared to the toric code, however, simplifies its implementation in physical quantum systems. The lattice structure of the planar surface codes are illustrated in Fig. 1.
@@ II.2.1 BP over GF(4)
The idea of maximum likelihood decoding in quantum error correction is to find the most likely error coset given the error syndrome. This can be expressed as [22]
@@
where the term $ES$ denotes the coset formed by the error $E$ and the stabilizer $S$. However, finding this maximum likelihood solution is computationally exponential [41]. Belief propagation decoding algorithm simplifies the above objective to
@@
where the probability of each individual error $E_{j}$ given the syndrome $s$ is maximized independently for $j=1$ to $n$. Belief propagation makes two simplifications to the objective, compromising its performance on highly degenerate quantum codes.
@@
Then, for each variable node, if all log-likelihood ratios for Pauli errors are greater than 0, then the hard decision result is I; otherwise, it is the Pauli error with the smallest LLR value. The hard decisions of all qubits combine to form the error estimate $\hat{E}$. If $\hat{E}$ corrects the error syndrome, the algorithm converges; otherwise, it returns to the horizontal update for a new iteration. If the maximum number of iterations is reached, decoding is considered to have failed.
BP can be performed under either parallel or serial scheduling [42]. Under parallel scheduling, all variable nodes and check nodes update their messages simultaneously at each iteration, allowing for hardware parallelization and constant time complexity. On the other hand, the serial scheduling updates the messages of variable nodes and check nodes sequentially. This method leads to faster convergence and slightly higher accuracy, though it has a more complex hardware implementation and higher time complexity.
@@ II.2.2 BP as gradient optimization
The target variables for gradient updates are the a posteriori estimates of the probability of each type of Pauli error occurring at each variable node. Specifically, for the $e\in\{X,Y,Z\}$ type Pauli error on the $n$-th variable node, the a posteriori probability update is
@@
where $\widetilde{m}_{c_{i}\rightarrow v_{j}}$ is the a posteriori probability summation term approximated using $\tanh$
@@ II.2.3 Hardness of BP in decoding surface codes
The decoding accuracy of BP in decoding surface codes is unsatisfactory due to several reasons, including cycles, degeneracy and trapping sets.
@@
High degeneracy leads to BP decoding objectives Eq. (2) deviating from true targets Eq. (1) [22, 37]. This problem becomes particularly severe when errors can combine with stabilizers to form error types that are symmetric to original ones. In such cases, BP’s marginal probability calculations for the both error types will be completely consistent. In surface codes, weights of the stabilizers are all no greater than 4, often much less than the weights of the logical operators. This symmetric degeneracy results in incorrect convergence of BP decoding.
@@ III Belief Propagation Optimization
Surface codes violate the assumptions of BP’s objective function and probability independence, making BP decoding difficult to converge. These issues are manifested in symmetric trapping sets, where the local structure of each node is identical, and message updates are symmetric under specific error patterns. Breaking the former symmetry involves modifying the code structure, while in this section, we focus on breaking the latter, the symmetry of message updates.
@@ III.1.1 Update smoothing
The Exponential Weighted Average (EWA) updates and smooths data series by combining the current observation with a weighted average of the previous value, expressed as
@@
In the field of machine learning, momentum methods utilize an idea similar to EWA, adding a momentum term to allow gradient descent to accumulate historical information
@@
Momentum-BP can be implemented under both parallel and serial scheduling. Here, we provide the latter in algorithm 1, where the initialization and check node update remain the same as in traditional BP. Our approach computes the a posteriori probability before performs the vertical update. The posterior update rule of Momentum-BP is shown as Eq. (16) and Eq. (17) where $g_{v_{j},e}^{(t)}$ are the accumulated momentum term and $Q_{v_{j},e}^{(t)}$ are the a posteriori probability for the $j$-th variable node regarding the Pauli error $e$ in the $t$-th iteration. Notice that we transformed Eq. (8), so that the accumulated term is the posterior value from the previous iteration. In the variable node update step, we subtract the original message sent from the check node rather than the momentum-smoothed message to ensure stability and prevent overcompensation.
@@ III.1.2 Adaptive step size
The oscillation of message updates in trapping sets can also be attributed to the step size. In traditional LLR-BP, all variable node posterior updates result from directly accumulating messages from neighboring check nodes, which can become symmetric and cause oscillations. From the gradient perspective, if the step size is too large, the optimization algorithm is likely to oscillate; if the step size is too small, the algorithm is unable to update effectively. In practical decoding, the update frequency of different variable nodes may be uneven, with some requiring more rapid updates while others need to remain stable.
@@
This method adaptively alters the step size by accumulating historical gradient information. For instance, when the estimate of a variable node for a certain Pauli error changes significantly, the algorithm reduces the step size to counteract oscillation. Eq. (21) takes effect after the first iteration to ensure initial update stability. Here, $\alpha$ is the boost learning rate, providing an update impetus as the algorithm begins to accumulate gradients. This learning rate is relatively insensitive; a moderate value suffices. In all our tests, $\eta$ is fixed at 5. While AdaGrad-BP may face issues with accumulated gradients causing the step size to approach zero after many iterations, this does not significantly impact the algorithm. In practice, AdaGrad can reduce the number of iterations, allowing most error patterns to converge before the gradients become too small.
Regrettably, combining Momentum-BP and AdaGrad-BP using the Adam optimizer or other methods has not yielded satisfactory results. On one hand, the Adam algorithm introduces additional hyperparameters whose optimal values for BP decoding cannot be directly borrowed from deep learning, thereby complicating the algorithm. On the other hand, Momentum and AdaGrad approach similar message update optimizations from different perspectives, and a straightforward combination of the two can lead to overcompensation.

### v3

@@ Abstract
Quantum error correction is crucial for universal fault-tolerant quantum computing. Highly accurate and low-time-complexity decoding algorithms play an indispensable role in ensuring quantum error correction works effectively. Among existing decoding algorithms, belief propagation (BP) is notable for its nearly linear time complexity and general applicability to stabilizer codes. However, BP’s decoding accuracy without post-processing is unsatisfactory in most situations. This article focuses on improving the decoding accuracy of BP over GF(4) for surface codes. We first propose Momentum-BP and AdaGrad-BP, inspired by machine learning optimization techniques, to reduce oscillation in message updating and break the symmetric trapping sets. We further propose EWAInit-BP, which adaptively updates initial probabilities and provides a 1 to 3 orders of magnitude improvement over traditional BP for planar surface code, toric code, and XZZX surface code without any post-processing method, showing high decoding accuracy even under parallel scheduling. The theoretical $O(1)$ time complexity under parallel scheduling and high accuracy of EWAInit-BP make it a promising candidate for high-precision real-time decoders. Meanwhile, the ideas of the Momentum-BP, AdaGrad-BP and EWAInit-BP provide promising approaches to improve the decoding accuracy of BP to get rid of its reliance on post-processing.
@@ I Introduction
Information in quantum computing devices is highly susceptible to noise, leading to information loss and computational errors [1]. To address this issue, quantum error correction (QEC) introduces redundant physical qubits and encodes the information in a certain subspace of the state space of all the physical qubits to protect it. So far at least, QEC is an essential step towards universal fault-tolerant quantum computation [2, 3, 4]. To ensure QEC works, the decoder of a QEC code should output an accurate and fast estimation of error according to the error syndromes.
@@
Several decoding algorithms have achieved notable success, including minimum-weight perfect matching (MWPM) [5] algorithm, union-find (UF) algorithm [6], maximum likelihood (ML) algorithm [7], matrix-product-states-based (MPS-based) algorithm [7], renormalization group algorithm [8], and neural-network-based (NN-based) algorithm [9, 10, 11, 12, 13, 14, 15, 16]. Among these, ML, MWPM, and MPS-based algorithms offer high accuracy but come with high complexity. In contrast, the UF algorithm provides near-linear complexity with slightly lower accuracy. NN-based decoders have to balance scalability, decoding accuracy, and time cost [13], and may struggle with generalizability across different codes. Consequently, existing decoding algorithms rarely meet the requirements for both high accuracy and low complexity in quantum error correction decoding [17, 18].
@@
Belief Propagation (BP) algorithm [19] are applicable to any stabilizer code and can achieve a complexity of $O(j\tau N)$ [20], where $j$ is the average weight of the rows and columns of the check matrix, and $\tau$ is the number of iterations. Additionally, parallel scheduling of BP can be efficiently implemented in hardware, theoretically achieving constant-time complexity [21]. However, traditional BP exhibits low decoding accuracy for many quantum error correcting codes and even cannot achieve the code capacity threshold for surface codes.
@@
BP can be implemented over both GF(4) [22, 23, 24, 25, 26, 20, 27] and GF(2) [28, 29, 30, 31], where the Pauli operators are represented. There have been some efforts improving decoding accuracy of BP over GF(4) and GF(2). Over GF(4), MBP [20] achieves a threshold of 14.5% to 16% for the planar surface code by modifying the step size in the a posteriori probability updates. However, MBP is limited to serial scheduling, and the outer loop for searching step sizes increases its time complexity. Over GF(2), GBP [31] achieves a 14% threshold by combining generalized belief propagation with multiple rounds of decoding. This algorithm can be implemented in parallel, but the number of outer loops increases with the code length, adding to its time complexity. BP-OTS [32] periodically detects and breaks trapping sets using manually set parameters, allowing parallel implementation while maintaining the original complexity. It outperforms MBP for short codes but does not achieve a significant threshold.
@@
This article employs three strategies to improve BP: optimizing message updates by optimization methods in machine learning, and optimizing a priori probabilities by utilizing a posteriori probabilities output by previous iterations. First, inspired by the similarity between BP message update and gradient descent of the energy function, we leverage optimization methods in machine learning to optimize the message update rules, and propose Momentum-BP and AdaGrad-BP. In addition to improving decoding accuracy, experiments on trapping sets in surface codes show that Momentum-BP and AdaGrad-BP can smooth the message update process and reduce oscillation amplitude.
Furthermore, realizing the importance of a priori and a posteriori probabilities involved in the iteration of BP, we propose EWAInit-BP which dynamically updates the a priori probabilities involved in each iteration by utilizing the a posteriori probabilities output by previous iterations. Experiments on short toric codes, planar surface codes, and XZZX surface codes (under biased noise) show that EWAInit-BP outperforms existing BP improvements without post-processing and outer loops, under both parallel and serial scheduling. Notably, the time complexity of all proposed improvements remains $O(j\tau N)$. Thus, EWAInit-BP can be implemented with theoretically constant complexity in parallel while maintaining decoding accuracy.
@@ II.1 Surface codes
A stabilizer code is defined by an abelian subgroup $\mathcal{S}$ of the Pauli group $\mathcal{P}_{n}$ excluding $-I$, and composed of operators $S\in\mathcal{P}_{n}$ such that $S|\psi\rangle=|\psi\rangle$ for all codewords $|\psi\rangle$ in the code space. The code space is the simultaneous $+1$ eigenspace of all elements of $\mathcal{S}$. The stabilizers must commute with each other, which means $[S_{i},S_{j}]=0$ for $\forall S_{i},S_{j}\in\mathcal{S}$. The dimension of the code space is determined by the number of independent generators of $\mathcal{S}$. For a stabilizer code with $N$ physical qubits and $K$ logical qubits, $\mathcal{S}$ has $N-K$ independent generators.
@@
The stabilizers are used to detect errors by measuring their eigenvalues. The set of measurement outcomes, known as the error syndrome, contains information about the errors on physical qubits. The logical operators of a stabilizer code are Pauli operators that commute with all elements of $\mathcal{S}$ but are not in $\mathcal{S}$, thus they are in $\mathcal{N(S)\backslash S}$ where $\mathcal{N(S)}=\{E:ES=SE,\ \forall S\in\mathcal{S}\}$. They act on the encoded logical qubits without disturbing the code space, thus they can lead to undetectable errors. The code distance $D$ is defined as the minimum weight of logical operators. A quantum stabilizer code can be represented by $[[N,K,D]]$.
@@
Surface codes may be the most notable stabilizer codes in recent years. Surface codes are a class of quantum Low-Density Parity-Check (LDPC) codes [39] where qubits are placed on a two-dimensional lattice, relying on the topology of the surface to protect quantum information [40]. Logical qubits are encoded in global degrees of freedom, and larger lattices provide larger code distances. There are some common properties of surface codes, such as the code distance scales with the code length as $O(\sqrt{N})$, while the code rate is asymptotically zero. In this article, toric codes refer to those with closed boundaries, and planar surface codes refer to those with open boundaries. The stabilizers of both types are either Pauli $X$ or Pauli $Z$. Besides, XZZX surface codes refer to the toric codes whose stabilizers are in the form of $XZZX$.
@@
A $[[2L^{2},2,L]]$ toric code has a 2-dimensional torus topology, and the closed boundaries makes it translationally invariant to the error syndrome, which is beneficial to decoding algorithms such as BP and neural networks. The Euler characteristic of the torus is 0, but the X-type and Z-type stabilizers of the toric code both have a redundancy, resulting in a total of two encoded logical qubits.
@@
A $[[2L^{2}-2L+1,1,L]]$ planar surface code is defined on a surface with open boundaries, which makes the stabilizers near the boundaries having lower weights. This reduction in stabilizer weight results in a slight decrease in the error correction capability compared to the toric code; however, it simplifies implementation in physical quantum systems.. The lattice structure of the planar surface codes are illustrated in Fig. 1.
@@ II.2.1 BP over GF(4)
The idea behind maximum likelihood decoding in quantum error correction is to find the most likely error coset given the error syndrome. This can be expressed as [22]
@@
where the term $ES$ denotes the coset formed by the error $E$ and the stabilizer $S$. However, finding this maximum likelihood solution is computationally expensive and exponential [41]. Belief propagation decoding algorithm simplifies the above objective to
@@
where the probability of each individual error $E_{j}$ given the syndrome $s$ is maximized independently for $j=1$ to $n$. Belief propagation makes two simplifications to the objective, which compromise its performance on highly degenerate quantum codes.
@@
Then, for each variable node, if all log-likelihood ratios for Pauli errors are greater than 0, then the hard decision result is the identity $I$; otherwise, it is the Pauli error with the smallest LLR value. The hard decisions of all qubits combine to form the error estimate $\hat{E}$. If $\hat{E}$ corrects the error syndrome, the algorithm converges; otherwise, it returns to the horizontal update for a new iteration. If the maximum number of iterations is reached, decoding is considered to have failed.
BP can be performed under either parallel or serial scheduling [42]. Under parallel scheduling, all variable nodes and check nodes update their messages simultaneously in each iteration, allowing for hardware parallelization and constant time complexity. On the other hand, the serial scheduling updates the messages of variable nodes and check nodes sequentially. This method leads to faster convergence and slightly higher accuracy, though it has a more complex hardware implementation and higher time complexity.
@@ II.2.2 BP as gradient optimization
The target variables for gradient updates are the a posteriori estimates of the probabilities of each type of Pauli error occurring at each variable node. Specifically, for the $e\in\{X,Y,Z\}$ type Pauli error on the $n$-th variable node, the a posteriori probability update is
@@
where $\widetilde{m}_{c_{i}\rightarrow v_{j}}$ is the a posteriori probability summation term, approximated using the $\tanh$ function
@@ II.2.3 Hardness of BP in decoding surface codes
The decoding accuracy of BP in decoding surface codes is unsatisfactory due to several reasons, including cycles, degeneracy, and trapping sets.
@@
High degeneracy leads to BP decoding objectives Eq. (2) deviating from true targets Eq. (1) [22, 37]. This problem becomes particularly severe when errors can combine with stabilizers to form error types that are symmetric to the original ones. In such cases, BP’s marginal probability calculations for the both error types will be completely consistent. In surface codes, the weights of the stabilizers are all no greater than 4, often much less than the weights of the logical operators. This symmetric degeneracy results in incorrect convergence of BP decoding.
@@ III Belief Propagation Optimization
Surface codes violate the assumptions of BP’s objective function and probability independence, making BP decoding difficult to converge. These issues manifest in symmetric trapping sets, where the local structure of each node is identical, and message updates are symmetric under specific error patterns. Breaking the former symmetry involves modifying the code structure, while in this section, we focus on breaking the latter, the symmetry of message updates.
@@ III.1.1 Update smoothing
The Exponential Weighted Average (EWA) method updates and smooths data series by combining the current observation with a weighted average of the previous value, expressed as
@@
In machine learning, momentum methods utilize an idea similar to EWA, adding a momentum term to accumulate historical information in gradient descent
@@
Momentum-BP can be implemented under both parallel and serial scheduling. Here, we present the latter in algorithm 1, where the initialization and check node update remain the same as in traditional BP. Our approach computes the a posteriori probability before performs the vertical update. The posterior update rule of Momentum-BP is shown as Eq. (16) and Eq. (17) where $g_{v_{j},e}^{(t)}$ are the accumulated momentum term and $Q_{v_{j},e}^{(t)}$ are the a posteriori probability for the $j$-th variable node regarding the Pauli error $e$ in the $t$-th iteration. Notice that we transformed Eq. (8), so that the accumulated term is the posterior value from the previous iteration. In the variable node update step, we subtract the original message sent from the check node rather than the momentum-smoothed message to ensure stability and prevent overcompensation.
@@ III.1.2 Adaptive step size
Oscillation of message updates in trapping sets can also be attributed to the step size. In traditional LLR-BP, all variable node posterior updates result from directly accumulating messages from neighboring check nodes, which can become symmetric and cause oscillations. From the gradient perspective, if the step size is too large, the optimization algorithm is likely to oscillate; if the step size is too small, the algorithm is likely unable to update effectively. In practical decoding, the update frequency of different variable nodes may be uneven, with some requiring more rapid updates while others need to remain stable.
@@
This method adaptively adjusts the step size by accumulating historical gradient information. For instance, when the estimate of a variable node for a certain Pauli error changes significantly, the algorithm reduces the step size to counteract oscillation. Eq. (21) takes effect after the first iteration to ensure initial update stability. Here, $\alpha$ is the boost learning rate, providing an update impetus as the algorithm begins to accumulate gradients. This learning rate is relatively insensitive; a moderate value suffices. In all our tests, $\eta$ is fixed at 5. While AdaGrad-BP may face issues with accumulated gradients causing the step size to approach zero after many iterations, this does not significantly impact the algorithm. In practice, AdaGrad can reduce the number of iterations, allowing most error patterns to converge before the gradients become too small.
Unfortunately, combining Momentum-BP and AdaGrad-BP using the Adam optimizer or other methods has not yielded satisfactory results. On one hand, the Adam algorithm introduces additional hyperparameters whose optimal values for BP decoding cannot be directly borrowed from deep learning, thereby complicating the algorithm. On the other hand, Momentum and AdaGrad approach similar message update optimizations from different perspectives, and a straightforward combination of the two can lead to overcompensation.